A teller has 49 $10 and $20 bills in her cash drawer. The value of the bills is $870. How many $10 bills does she have?

Answer by scott8148(6628)

let x=tens, so 49-x=twenties ... 10(x)+20(49-x)=870 ... 10x+980-20x=870